# Inferring Resource-Oriented Intentions using LLMs for Static Resource Leak Detection

### v3

@@ V-A Effectiveness in Resource Leak Detection (RQ1)
In this evaluation, we assess the effectiveness of InferROI in detecting resource leaks using two datasets.
@@ V-A2 Baselines
We compare InferROI with three well-maintained industrial static detectors, as evaluated in previous studies [8, 13].
@@ V-A3 Procedure
Following previous study [8], we adopt the metrics Bug Detection Rate and False Alarm Rate to evaluate the detection effectiveness of InferROI and baseline detectors.
TABLE I: Resource Leak Detection on DroidLeaks Dataset
Detector # Detected Bugs (Bug Detection Rate) # False Alarms (False Alarm Rate)
SpotBugs 6 (6.9%) 0 (0%)
Infer 37 (43.0%) 16 (18.6%)
PMD 10 (11.6%) 6 (7.0%)
InferROI (ours) 51 (59.3%) 16 (18.6%)
TABLE II: Resource Leak Detection on JLeaks Dataset
Detector # Detected Bugs (Bug Detection Rate)
SpotBugs* 44 (24.3%)
Infer* 5 (9.4%)
PMD 341 (43.5%)
InferROI (ours) 508 (64.8%)
    Infer and SpotBugs encounter compilation issues with the majority of test instances.
For each bug in our experiments, we apply InferROI to its buggy version and fixed version. If InferROI reports a resource leak in the buggy version related to the corresponding concerned resource type, it is considered to have correctly detected the bug. Conversely, if InferROI reports a resource leak in the fixed version for the corresponding concerned resource type, it signifies a false alarm. The two metrics are calculated as follows:
Bug Detection Rate=# detected bugs in buggy versions# total experimented $\textit{Bug Detection Rate}=\frac{\text{\# detected bugs in buggy versions}}{\text{\# total experimented bugs}}$
False Alarm Rate=# reported false alarms in fixed versions# total experimented $\textit{False Alarm Rate}=\frac{\text{\# reported false alarms in fixed versions}}{\text{\# total experimented bugs}}$
@@ V-A4 Results
DroidLeaks: InferROI demonstrates a bug detection rate of 59.3% with a false alarm rate of 18.6% across 86 bugs in DroidLeaks. Compared to the baseline detectors, InferROI achieves notably higher bug detection, detecting 14~45 more bugs, while maintaining a comparable false alarm rate.
JLeaks: InferROI showcases a bug detection rate of 64.8% with a false alarm rate of 24.0% across 784 bugs in JLeaks. In contrast to the baseline detectors, InferROI achieves significantly superior bug detection, identifying 167~503 more bugs in JLeaks. It’s important to note that we do not report the false alarm rate of the baselines due to the unavailability of data necessary for its calculation in current JLeaks dataset.
@@
Analysis: (i) Bug Detection Analysis: The higher bug detection rate of InferROI can be attributed to the advantages of the resource-oriented intention inference utilized by InferROI, as discussed in the insights provided in Section II (see RQ3.a for the resource coverage assessment). However, it is essential to acknowledge that the resource-oriented intention inference cannot entirely replace the value of more sound program analysis techniques. For instance, in the case of buggy versions not detected by InferROI, a substantial portion is caused by the intricacies of Android’s lifecycle management and callback mechanisms. This observation indicates an important future direction, i.e., to explore better integration of LLM-based resource-intention inference with more advanced program analysis techniques, which could enable more general and powerful resource leak detection. (ii) False Alarm Analysis: False alarms reported by InferROI primarily stem from incorrect or missing intentions. When InferROI incorrectly identifies an acquisition intention or overlooks a release intention, a false alarm occurs. To address this issue, we intend to explore the integration of LLM-based prompt engineering techniques, such as in-context learning and chain-of-thought [27], to enhance the capabilities of LLMs. Notably, a significant portion of these “false alarms” arises from misclassifications due to the quality of the ground truth. Since both datasets are curated based on bug-fixing commits, certain “fixed” versions still contain resource leaks. This underscores the ongoing challenge of ensuring the quality of resource leak datasets.
@@
Summary: InferROI demonstrates a bug detection rate of 59.3% with a false alarm rate of 18.6% across 86 bugs in DroidLeaks, and a bug detection rate of 64.8% with a false alarm rate of 24.0% across 784 bugs in JLeaks. This performance shows a promising potential of detecting resource leaks based on resource-oriented intentions.
@@ V-B Effectiveness in Open-Source Project Scanning (RQ2)
We apply InferROI on real-world Java projects to evaluate its efficacy in detecting previously-unknown resource leaks.
@@ V-B1 Data
We crawl 115 Java open-source projects with more than 50 stars, which are created after December 31, 2021. The date filtering condition is used to avoid the projects are included in the training data of the employed GPT-4, which is trained using a huge amount of data up to September 2021. To save time and the cost of calling the GPT-4 API (https://platform.openai.com/docs/api-reference), we randomly filter 100 methods from 13 projects, by matching 20 common resource terms reported in [28], instead of completely scaning these projects. The employed terms are listed as follows: stream, reader, client, writer, lock, player, connection, monitor, gzip, ftp, semaphore, mutex, stream, camera, jar, buffer, latch, socket, database, scanner, cursor.
@@ V-B2 Procedure
We apply InferROI on these 100 methods to detect resource leaks. For each reported leak, we first manually annotate whether it is a true bug by reading code and consulting related information. For the true bugs, we then submit the corresponding bug-fix pull requests and help the project developers review them.
@@ V-B3 Results
In the 100 methods, InferROI reports 29 resource leaks and 26 are annotated as true bugs. At the submission time, 7 of the true bugs have been confirmed by the project developers, and the pull requests have been merged.
@@
Comparison with Static Detector: The static detector, PMD, reports 20 resource leaks in the 100 methods, with 18 of them being true bugs. Among these 18 bugs identified by PMD, InferROI successfully detects 17 of them. The lone bug missed by InferROI involves a method comprising over 400 lines of code, posing challenges for LLM processing. Conversely, InferROI identifies 9 additional true bugs not caught by PMD. Noteworthy among these are resource types such as ManagedBuffer from the Apache Spark library, InMemoryDirectoryServer from UnboundID LDAP, and a custom type JDBCConnection specific to the project.
Comparison with Resource Mining Approach: Of the 26 bugs, 3 are beyond the scope of the state-of-the-art resource acquisition/release API pair (RAR pair) mining approach MiROK [28]. MiROK primarily concentrates on extracting the definitions of abstract RAR pairs, which differs from our objective of directly detecting resource-oriented intentions in the usages of RAR pairs. Nonetheless, the abstract RAR pairs mined by MiROK can aid in filtering suspicious methods for InferROI, similar to the 20 resource terms employed in this evaluation.
@@
Summary: InferROI successfully detects 26 previously-unknown resource leak bugs in 100 methods from real-world open-source projects. The results demonstrate the practical capability of InferROI in effectively detecting resource leaks in real-world software projects.
V-C Accuracy in Resource-Oriented Intention Inference (RQ3.a)
@@
V-C1 Procedure
To evaluate the effectiveness of InferROI’s intention inference, two authors independently annotate all three types of resource-oriented intentions present in the 172 code snippets collected from the DroidLeaks dataset. In cases of disagreement, they engage in discussions to reach a final decision, ensuring consistent and accurate annotations. These manual annotations serve as the ground truth intentions ($GTs$) for the evaluation.
Subsequently, we apply InferROI to the code snippets to infer the intentions (denoted as $Preds$), and then calculate the precision and recall of the inference based on the manual annotations. Precision represents the proportion of intentions inferred by InferROI that align with the ground truth, while recall indicates the proportion of true intentions from the ground truth that InferROI correctly infers. The precision and recall calculations are as follows:
@@ V-C1 Procedure
Furthermore, we evaluate the resource coverage of InferROI for DroidLeaks and JLeaks and compare it with PMD. For InferROI, we determine the portion of the 28 (DroidLeaks) and 368 (JLeaks) resource types covered by the inferred intentions.
V-C2 Results
InferROI achieves a precision of 74.6% and a recall of 81.8% in resource-oriented intention inference, indicating its ability to relatively accurately and comprehensively identify intentions within the code snippets. Moreover, InferROI demonstrates superior versatility in addressing a broader spectrum of resource types, attaining a coverage of 67.9% (19 types) for DroidLeaks and 60.1% (221 types) for JLeaks, compared to PMD’s coverage of 32.1% (9 types) and 12.8% (47 types), respectively.
@@ V-C2 Results
V-D Ablation Study (RQ3.b)
We conduct an ablation study to analyze and compare the contributions of two key components of InferROI: the LLM-based resource-oriented intention inference and the static resource leak detection.
V-D1 Procedure
To perform the ablation study, we create two prompt templates to enable GPT-4 to directly detect resource leaks in code, resulting in two resource leak detectors: GPTLeak and GPTLeak-chain. The prompt templates for GPTLeak and GPTLeak-chain are illustrated in Figure 5, and they are structured similarly to the template utilized by InferROI. For GPTLeak, the instruction is straightforward, while for GPTLeak-chain, an additional detection instruction (the last sentence) is incorporated into the template used by InferROI to provide specific guidance to the GPT-4 model during the decision-making process. We apply both GPTLeak and GPTLeak-chain detectors to the 172 code snippets collected from the DroidLeaks dataset. Subsequently, we calculate the bug detection rate and false alarm rate for each detector.
Figure 5: Prmopt Templates for the Two GPT-based Detectors
V-D2 Results
As shown in Table III, InferROI substantially outperforms both GPTLeak and GPTLeak-chain, achieving the highest bug detection rate and the lowest false alarm rate. These findings show the effectiveness of InferROI’s combination of the LLM-based intention inference and the static analysis in detecting resource leaks. While LLMs offer extensive knowledge and understanding of code, they may still lack robust reasoning capacity to accurately detect resource leaks. On the other hand, the static analysis also provides a more rigorous approach, which can address the limitations of LLMs and reduce the potential for false positives and false negatives in the detection process. By synergistically combining the strengths of both approaches, InferROI achieves a balanced and effective resource leak detection mechanism, ultimately enhancing the overall accuracy and reliability of its results.
An interesting observation is that GPTLeak-chain performs worse than GPTLeak when utilizing the instruction chain for the intention inference. This phenomenon may suggest that the LLMs detect resource leaks using different logic than the procedure of static analysis-based detection. Explicitly inferring the three types of resource-oriented intentions proves to be beneficial for boosting static resource leak detection, but appears to be unnecessary for LLMs. This observation emphasizes the significance of prompt design when applying LLMs for resource leak detection.
TABLE III: Ablation Study Results
Detector # Detected Bugs (Bug Detection Rate) # False Alarms (False Alarm Rate)
GPTLeak 44 (51.2%) 18 (20.9%)
GPTLeak-chain 32 (37.2%) 36 (41.9%)
InferROI (ours) 51 (59.3%) 16 (18.6%)

### v4

@@ V-A Effectiveness in Resource Leak Detection (RQ1)
In this evaluation, we assess the effectiveness of InferROI with GPT-4 in detecting resource leaks using two datasets.
@@ V-A2 Baselines
In RQ1, we compare InferROI with three well-maintained industrial static detectors, as they have been evaluated in both DroidLeaks and JLeaks papers [8, 13].
@@ V-A3 Procedure
For each bug in our experiments, we apply InferROI to both its buggy version and its fixed version. If InferROI reports a resource leak in the buggy version related to the concerned resource type, it is considered a true positive (TP); otherwise, there is a false negative (FN). Conversely, if InferROI reports a resource leak in the fixed version for the corresponding resource type, it is a false positive (FP). Based on this, we count the numbers of TPs, FPs, and FNs, and calculate Bug Detection Rate (BDR) and False Alarm Rate (FAR) following a previous study [8].
$BDR=\frac{\text{\# TP}}{\text{\# Buggy Versions}},\quad FAR=\frac{\text{\#FP}}{\text{\# Fixed Versions}}$
Additionally, to better reflect the overall effectiveness of InferROI, we calculate the $F1$ score based on Precision and Recall:
$F1=\frac{2\cdot\textit{Precision}\cdot\textit{Recall}}{\textit{Precision}+\textit{Recall}},$
where Precision and Recall are calculated by #$\nicefrac{{\text{\#TP}}}{{(\text{\#TP}+\text{\#FP})}}$ and #$\nicefrac{{\text{\#TP}}}{{(\text{\#TP}+\text{\#FN})}}$, respectively. Note that we do not report Precision and Recall separately in the experimental results, as their implications are effectively captured by Bug Detection Rate and False Alarm Rate.
@@ V-A4 Results
TABLE I: Resource Leak Detection on DroidLeaks Dataset
  Detector #TP #FP BDR $\uparrow$ FAR $\downarrow$ F1 $\uparrow$
SpotBugs 6 0 6.9% 0% 0.129
Infer 37 16 43.0% 18.6% 0.532
PMD 10 6 11.6% 7.0% 0.196
InferROI 51 16 59.3% 18.6% 0.667
TABLE II: Resource Leak Detection on JLeaks Dataset
  Detector #TP #FP BDR $\uparrow$ FAR $\downarrow$ F1 $\uparrow$
SpotBugs* 44 24.3%
Infer* 5 9.4%
PMD 341 40 43.5% 5.1% 0.585
InferROI 490 153 62.5% 19.5% 0.687
    Due to encountering compilation issues, SpotBugs and Infer are able to process only 181 and 53 instances, respectively.
@@
DroidLeaks: InferROI demonstrates a bug detection rate of 59.3% with a false alarm rate of 18.6% across 86 bugs in DroidLeaks. Compared to the baseline detectors, InferROI achieves notably higher bug detection, detecting 14~45 more bugs, while maintaining a comparable false alarm rate. Considering the overall effectiveness, InferROI achieves an $F1$ score of 0.667, significantly outperforming the baselines by 25.4% to 417.1%.
JLeaks: InferROI shows a bug detection rate of 62.5% with a false alarm rate of 19.5% across 784 bugs in JLeaks. Compared to the baseline detectors, InferROI achieves significantly superior bug detection, identifying 149~485 more bugs in JLeaks. As JLeaks does not release complete compilable project-level environments, we cannot run the baselines SpotBugs and INFER (which both require compiled projects) and Table II presents their bug detection rate as reported in JLeaks paper. For the baseline PMD which does not require compiled projects, we run PMD on JLeaks and report both its false alarm rate and bug detection rate. As shown in Table II, although PMD achieves lower false alarm rate than InferROI, InferROI achieves much better trade-off between bug detection rate and false alarm rate than PMD, i.e., InferROI achieves an F1 score of 0.687, outperforming PMD by 17.4%.
@@
Analysis: (i) Bug Detection Analysis: The higher bug detection rate of InferROI can be attributed to the advantages of the resource-oriented intention inference utilized by InferROI, as discussed in the insights provided in Section II (see RQ3.a for the resource coverage assessment). However, it is essential to acknowledge that the resource-oriented intention inference cannot entirely replace the value of more sound program analysis techniques. For instance, in the case of buggy versions not detected by InferROI, a substantial portion is caused by the intricacies of Android’s lifecycle management and callback mechanisms. This observation indicates an important future direction, i.e., to explore better integration of LLM-based resource-intention inference with more advanced program analysis techniques, which could enable more general and powerful resource leak detection. (ii) False Alarm Analysis: False alarms reported by InferROI primarily stem from incorrect or missing intentions. When InferROI incorrectly identifies an acquisition intention or overlooks a release intention, a false alarm occurs. To address this issue, we intend to explore the integration of LLM-based prompt engineering techniques, such as in-context learning, to enhance the capabilities of LLMs. Notably, a significant portion of these “false alarms” arise from incorrect ground truths. Since both datasets are curated based on bug-fixing commits, certain “fixed” versions still contain resource leaks. This underscores the ongoing challenge of ensuring the quality of resource leak datasets.
@@
Summary: InferROI demonstrates a bug detection rate of 59.3% with a false alarm rate of 18.6% across 86 bugs in DroidLeaks, and a bug detection rate of 62.5% with a false alarm rate of 19.5% across 784 bugs in JLeaks. Considering overall effectiveness, InferROI significantly outperforms the baselines on both datasets, with $F1$ scores of 0.667 and 0.687. This shows a promising potential of InferROI in detecting resource leaks.
@@ V-B Effectiveness in Open-Source Project Scanning (RQ2)
We apply InferROI with GPT-4 on real-world Java projects to evaluate its efficacy in detecting previously-unknown resource leaks.
@@ V-B1 Data
To save time and reduce the cost of querying the GPT-4 API, we employ two strategies to sample two set of methods from open-source projects.
Suspicious Methods Filtered by Matching Resources. We crawl 115 Java open-source projects with more than 50 stars, which are created after December 31, 2021. We randomly filter 100 suspicious methods from 13 projects, after matching 20 common resource terms reported in previous work [31], instead of completely scanning these projects. The employed terms are listed as follows: stream, reader, client, writer, lock, player, connection, monitor, gzip, ftp, semaphore, mutex, camera, jar, buffer, latch, socket, database, scanner, cursor.
Random Methods Sampled from Apache Lucene. To further evaluate the effectiveness of InferROI in detecting resource leaks in the scenarios without relying on matching common resource terms, we also randomly sample 100 methods from the latest version of Apache Lucene [32], a widely used text retrieval engine.
Interprocedural Scanning with CodeQL on Apache Lucene. To investigate the potential of leveraging inferred intentions to enhance interprocedural scanning, we integrate the Resource Acquisition and Resource Release intentions inferred by InferROI for Apache Lucene into CodeQL QL scripts. This integration is straightforward, as CodeQL supports the addition of custom resource types and their associated acquisition and release APIs. However, we have not yet incorporated Resource Reachability Validation, as modifying CodeQL’s flow analysis to support it presents significantly greater complexity.
@@ V-B2 Procedure
We apply InferROI on these 200 methods to detect resource leaks. For each reported leak, we first manually annotate whether it is a true bug by reading code and consulting related information. For the true bugs, we then submit the corresponding bug-fix pull requests and help the project developers review them. We run PMD on these methods and manually check the reported bugs. In RQ2, we use PMD instead of SpotBugs and Infer as our baseline, because (i) PMD shows the best detection rate on the large-scale benchmark JLeaks in RQ1, and (ii) PMD does not require compiled projects (while each instance in RQ2 is a method snippet).
TABLE III: Results of Open Source Scanning
  Detector 100 Suspicious Methods 100 Lucene Methods
#TP #FP #TP #FP
InferROI 26 (7 Accepted PRs) 3 3 0
PMD 18 2 0 0
@@ V-B3 Results
In the 100 suspicious methods filtered by matching resources, InferROI reports 29 resource leaks and 26 are annotated as true bugs. At the submission time, 7 of the true bugs have been confirmed by the project developers, and the pull requests have been accepted. PMD reports 20 resource leaks in the 100 methods, with 18 of them being true bugs. Among these 18 bugs identified by PMD, InferROI successfully detects 17 of them. The lone bug missed by InferROI involves a method comprising over 400 lines of code, posing challenges for LLM processing. Conversely, InferROI identifies 9 additional true bugs not caught by PMD. Noteworthy among these are resource types such as ManagedBuffer from the Apache Spark library, InMemoryDirectoryServer from UnboundID LDAP, and a custom type JDBCConnection specific to the project.
In the 100 methods sampled from Apache Lucene, InferROI reports 3 resource leaks, all of which are confirmed as true bugs, while PMD reports zero bugs. The 3 bugs detected by InferROI involve project-defined resource classes IndexReader, DirectoryReader, and DirectoryTaxonomyReader. These resources are acquired and released by calling the incRef and decRef methods, respectively, which are challenging for static detectors to address.
For the interprocedural scanning, CodeQL reports 34 leaks before integrating InferROI-inferred intentions and 49 leaks after integration. Among these, 29 (85.3%) and 41 (83.7%) are confirmed as true bugs, respectively.
@@
Summary: InferROI successfully detects 26 previously unknown bugs in 100 suspicious methods and 3 previously unknown bugs in 100 Lucene methods. The results demonstrate the practical capability of InferROI in effectively detecting resource leaks in real-world software projects. Integrating InferROI-inferred intentions enables CodeQL to identify 12 additional true bugs during interprocedural scanning on Apache Lucene.
V-C Ablation Study
We compare InferROI with three end-to-end GPT-based detectors that rely solely on prompting engineering to evaluate the contribution of static analysis.
V-C1 Procedure
The three end-to-end GPT-based detectors without static analysis are denoted as GPTLeak variants (The used prompts can be found in our replication package [33]).
  • GPTLeak employs a straightforward instruction to directly identify resource leaks in the provided code.
  • GPTLeak-exp employs an additional chain-of-thought (CoT) instruction “First, explain the behavior of the code” to explain code behavior, which have been proven effective in static bug detection in previous study [34].
  • GPTLeak-roi employs additional CoT instructions to infer resource-oriented intentions (ROIs) in code, similar to InferROI, before identifying resource leaks.
We apply the three detectors to the 172 code snippets collected from the DroidLeaks dataset and calculate the BDR, FAR, and $F1$ score for each detector.
V-C2 Results
As shown in Table  IV, InferROI outperforms three GPTLeak variants (including two CoT variants) in overall detection effectiveness, achieving a higher F1 score with comparable BDR and much lower FAR. In particular, the two CoT variants outperform the straightforward GPTLeak, indicating that the suitable design of CoT can help LLMs better identify bugs. In addition, InferROI further outperforms the two CoT variants, indicating the benefits of combining LLMs and static analysis, which is also the main insight of InferROI. More detailed analyses are presented below.
Comparison among CoT variants and the basic GPTLeak. Compared to the straightforward GPTLeak, the two CoT strategies in GPTLeak-exp and GPTLeak-roi show improvements in reducing false alarms while slightly decreasing the bug detection rate. For example, GPTLeak-roi reduces more than half of the false alarms compared to GPTLeak by employing CoT instructions to first inferring resource-oriented intentions (i.e., the same as InferROI). It actually confirms that our proposed intention inference is also beneficial for purely LLM-based resource leak detection. Overall, incorporating suitable CoT prompt is beneficial when applying LLMs for resource leak detection.
Comparison between InferROI and CoT variants. Although the two CoT variants (i.e., GPTLeak-exp and GPTLeak-roi) outperform the basic GPTLeak, InferROI further outperforms the two CoT variants in overall detection effectiveness. In particular, the static analysis can equip LLMs with more rigorous analysis, which can help reduce false positives in end-to-end LLM-based detection; while LLMs are also complementary to the static analysis as LLMs exhibit more general capabilities in code comprehension. As a result, by combining static analysis and LLMs, InferROI can achieve a balanced and effective resource leak detection mechanism, ultimately enhancing its overall accuracy and reliability.
TABLE IV: Ablation Study Results
  Detector #TP #FP BDR $\uparrow$ FAR $\downarrow$ F1 $\uparrow$
GPTLeak 57 58 66.3% 67.4% 0.567
GPTLeak-exp 52 37 60.5% 43.0% 0.594
GPTLeak-roi 49 26 57.0% 30.2% 0.609
InferROI 51 16 59.3% 18.6% 0.667
V-D Generalizability with Different LLMs (RQ4)
We conduct an experiment to investigate the generalizability of InferROI when integrated with different LLMs.
V-D1 Selection of LLMs
In addition to the advanced closed-source GPT-4, we explore the effectiveness and efficiency of InferROI when integrated with the following two state-of-the-art open-source LLMs.
  • Llama3-8B [35]: Meta’s Llama 3 is a family of large language models available in various sizes, featuring both pretrained and instruction-tuned generative text models. We use Llama3-8B, which has 8 billion parameters.
  • Gemma2-9B [36]: Google’s Gemma 2 models are lightweight, instruction-tuned large language models, built using the same technology as the Gemini models [37]. We employ the Gemma2-9B version, which has 9 billion parameters.
V-D2 Results
Table V presents the comparison in the effectiveness and efficiency of InferROI when integrated with different LLMs. Overall, INFERROI with different LLMs consistently outperforms existing baselines (i.e., SpotBugs, Infer, and PMD) in F1 score. The results show that open-source Gemma2-9B achieves comparable effectiveness (0.644 vs. 0.667 in $F1$ score) with GPT-4, while Llama3-8B shows relatively lower metrics. The two smaller open-source models, Llama3-8B and Gemma2-9B, have longer response times than the larger GPT-4 due to the lack of optimizations like batch processing.
Summary: InferROI demonstrates substantial generalizability when integrated with different LLMs, including closed-source GPT-4 and open-source LLMs, maintaining relatively consistent effectiveness in resource leak detection.
TABLE V: Results of Different LLMs on DroidLeaks Dataset
  LLM #TP #FP BDR $\uparrow$ FAR $\downarrow$ F1 $\uparrow$ Time
Llama3-8B 39 18 45.3% 20.9% 0.545 7.1s
Gemma2-9B 48 15 55.8% 17.4% 0.644 11.4s
GPT-4 51 16 59.3% 18.6% 0.667 8.7s
V-E Accuracy in Resource-Oriented Intention Inference (RQ5)
@@
V-E1 Procedure
To evaluate the effectiveness of InferROI’s intention inference, two authors independently annotate all three types of resource-oriented intentions present in the 172 code snippets collected from the DroidLeaks dataset. In cases of disagreement, they engage in discussions to reach a final decision, ensuring consistent and accurate annotations. These manual annotations serve as the ground truth intentions ($GTs$) for the evaluation. Subsequently, we apply InferROI to the code snippets to infer the intentions (denoted as $Preds$), and then calculate the precision and recall of the inference based on the manual annotations as follows:
@@ V-E1 Procedure
We compare the intention inference of InferROI with MiROK [31], which is a state-of-the-art technique for mining RAR (Resource Acquisition and Release) API pairs. Specifically, we use the abstract RARs reported by MiROK to match APIs in the code for identifying intentions and then calculate precision and recall using the aforementioned equations. Note that MiROK does not support VAL intentions, so its metrics are calculated only for ACQ and REL intentions. Furthermore, we assess the coverage of InferROI-inferred intentions for the 28 resource types in DroidLeaks and the 368 resource types in JLeaks, and then compare this coverage with that of PMD.
V-E2 Results
InferROI achieves a precision of 74.6% and a recall of 81.8% in resource-oriented intention inference, indicating its ability to relatively accurately and comprehensively identify intentions within the code snippets. In contrast, intention identification based on MiROK yields only 54.0% precision and 42.3% recall, which is significantly lower than InferROI. Moreover, InferROI demonstrates superior versatility in addressing a broader spectrum of resource types, attaining a coverage of 67.9% (19 types) for DroidLeaks and 60.1% (221 types) for JLeaks, compared to PMD’s coverage of 32.1% (9 types) and 12.8% (47 types), respectively.